# Complete non-ambiguous trees and associated permutations: new enumerative results

### v2

@@ 2.3. Permutations, permutation patterns, and permutation graphs
For $n\geq 1$, we let $S_{n}$ be the set of permutations of size $n$ (called $n$-permutations for short). We will usually use standard one-line notation for permutations. That is, a permutation $\pi\in S_{n}$ is a word $\pi=\pi_{1}\cdots\pi_{n}$ on the alphabet $[n]$ such that every letter in $[n]$ appears exactly once in $\pi$. By convention, the unique 00-permutation is simply the empty word. For $i,j\in[n],i\neq j$, we write $i\prec_{\pi}j$ if $i$ appears before (to the left of) $j$ in the permutation (word) $\pi$.
Another useful way of viewing permutations is their plot. In this view, an $n$-permutation is a set of dots in an $n\times n$ grid such that every row and every column of the grid contains exactly one dot. For a permutation $\pi=\pi_{1}\cdots\pi_{n}$, we construct its plot by putting dots in column $i$ and row $\pi_{i}$ for $i=1,\cdots,n$. When connecting permutations to CNATs, it will be convenient to label columns from left-to-right and rows from top-to-bottom in this representation, as in Figure 2 below.
111122223333444455556666111122223333444455556666
@@
A permutation $\pi=\pi_{1}\cdots\pi_{n}$ is said to be reducible if there exists $1\leq k<n$ such that $\pi_{1}\cdots\pi_{k}$ is a $k$-permutation. A permutation is irreducible if it is not reducible. Such an index $k$ is called a breakpoint of the permutation $\pi$. For example, the permutation $\textbf{312}564$ is reducible, while $561243$ is irreducible.111 The terminology of (in)decomposable, or sum (in)decomposable, is also used sometimes in the literature instead of (ir)reducible. A fixed point of a permutation $\pi$ is an index $j\in[n]$ such that $\pi_{j}=j$. A left-to-right minimum is a letter $\pi_{i}$ such that $\pi_{j}>\pi_{i}$ for all $j<i$. For example, the permutation $\textbf{5}6\textbf{1}243$ has two left-to-right minima $5$ and $1$.
[…]
111122223333444455556666111122223333444455556666333344441111222255556666
@@
This is a classical result on permutation graphs, see e.g. [19, Lemma 3.2]. From now on, we assume that all permutations are irreducible unless explicitly stated otherwise. One permutation that will play an important role in our paper, particularly in Section 3 and Theorem 4.24, is the decreasing permutation, defined by $\mathrm{dec}_{[}n]:=n(n-1)\cdots 1$ (for some $n\geq 1$). In $\mathrm{dec}_{[}n]$ every pair is an inversion, so the corresponding graph is the complete graph $K_{n}$.
Let $\pi$ be an $n$-permutation, and $\tau$ a $k$-permutation for some $k\leq n$. We say that $\pi$ contains the pattern $\tau$ if there exist indices $i_{1}<i_{2}<\cdots<i_{k}$ such that $\pi_{i_{1}},\pi_{i_{2}},\cdots,\pi_{i_{k}}$ appear in the same relative order as $\tau$. In other words, if we delete all columns and rows other than $(i_{1},\pi_{i_{1}}),\cdots(i_{k},\pi_{i_{k}})$ from the plot of $\pi$, we get the plot of $\tau$. In this case, we call $\pi_{i_{1}}\cdots\pi_{i_{k}}$ an occurrence of the pattern $\tau$. If $\pi$ does not contain the pattern $\tau$, we say that $\pi$ avoids $\tau$. For example, the permutation $\pi=561243$ contains two occurrences of the pattern $321$, given by the bolded letters: $\textbf{5}612\textbf{43}$, $5\textbf{6}12\textbf{43}$. However, it avoids the pattern $4321$ since there is no sequence of $4$ letters in decreasing order.
@@
Patterns of a permutation $\pi$ are linked to induced subgraphs of its permutation graph $G_{\pi}$ via the following observation. Let $\pi$ be an $n$-permutation, and $\tau$ a $k$-permutation, for some $k\leq n$. Suppose that $\pi_{1},\cdots,\pi_{k}$ is an occurrence of the pattern $\tau$ in $\pi$. Then the induced subgraph $G\left[\{\pi_{1},\cdots,\pi_{k}\}\right]$ is isomorphic to the permutation graph $G_{\tau}$. Conversely, if there exist vertices $\pi_{1},\cdots,\pi_{k}$ such that $G\left[\{\pi_{1},\cdots,\pi_{k}\}\right]$ is isomorphic to $G_{\tau}$, then $\pi_{1},\cdots,\pi_{k}$ is an occurrence of some pattern $\tau^{\prime}$ in $\pi$, such that the graphs $G_{\tau}$ and $G_{\tau^{\prime}}$ are isomorphic. This yields the following.
@@ 2.4. Complete non-ambiguous trees and associated permutations
Following tree terminology, for a NAT $T$, we call the dot lying in the top-left cell the root dot, or simply the root, of $T$. Similarly, a leaf dot is a dot with no dots below it in the same column or to its right in the same row. An internal dot is a dot which is not a leaf dot (this includes the root, unless the NAT is a single dotted cell). Given a NAT, it will be convenient to label the columns $1,2,\cdots n$ from left to right, and the rows $1,2,\cdots,m$ from top to bottom.
@@
It is straightforward to check that in a CNAT, every row and every column must have exactly one leaf dot (see e.g. [13, Section 4.3] or [11, Section 2.1]). The leaf dot in a column, resp. row, is simply its bottom-most, resp. right-most, dot. As such, the set of leaf dots of a CNAT $T$ of size $n$ forms the plot of an $n$-permutation $\pi$, which must be irreducible (see e.g. [11, Theorem 3.3]). We say that $\pi$ is the permutation associated with the CNAT $T$. For example, the CNAT on the left of Figure 4 has associated permutation $\pi=45312$. For a permutation $\pi$, we define the set $\mathrm{CNAT}\left([\right)\pi]$ to be the set of CNATs whose associated permutation is $\pi$, and $\mathrm{cnat}\left([\right)\pi]:=|\mathrm{CNAT}\left([\right)\pi]|$ to be the number of such CNATs. We say that a row or column of a CNAT is empty if it contains no dots other than its leaf dot.
2.5. Connecting CNATs to the Tutte polynomial
We are now ready to state the connection between CNATs and the Tutte polynomial. This is essentially a re-phrasing of [13, Proposition 32] if one uses the definition of Tutte polynomials with spanning tree activities (see [14, Definition 2.16]). Recall that for a graph $G$, we define $t_{G}:=T_{G}(1,0)$.
@@ 2.5. Connecting CNATs to the Tutte polynomial
$\mathrm{cnat}\left([\right)\pi]=t_{G_{\pi}}.$
@@ 3. Upper-diagonal CNATs and permutations
In [11], the authors showed that for the decreasing permutation $\mathrm{dec}_{[}n]=n(n-1)\cdots 1$, we have $\mathrm{cnat}\left([\right){\mathrm{dec}_{[}n]}]=(n-1)!$. Their proof is via so-called tiered trees, with the authors exhibiting a bijection between upper-diagonal fully tiered trees on $n$ vertices and permutations of $[n-1]$. In this section, we give a new proof of this result by exhibiting a direct bijection between the set $\mathrm{CNAT}\left([\right){\mathrm{dec}_{[}n]}]$ and the set of permutations of $[n-1]$. Our bijection also maps certain statistics of CNATs to statistics on the corresponding permutation. Following notation from [11], we call a CNAT upper-diagonal if its associated permutation is a decreasing permutation.
@@
A labelled CNAT is a pair $\mathcal{T}=(T,\lambda)$ where $T$ is an upper-diagonal CNAT of size $(n+1)$ for some $n\geq 0$ and $\lambda=\{\lambda_{1},\lambda_{2},\cdots,\lambda_{n}\}\subset\mathbb{N}^{n}$ is a set of $n$ distinct natural numbers. We call $\lambda$ the labels of $\mathcal{T}$, and say that $\mathcal{T}$ is a $\lambda$-labelled CNAT.
In the above, we think of $\lambda$ as labelling the columns of the CNAT $T$ other than the right-most column, as in Figures 5 and 6 below (the labels are written on top of the columns). With a slight abuse of notation, we assume that $\lambda_{1}<\lambda_{2}<\cdots<\lambda_{n}$, and use $\lambda$ to denote both the set of labels and the ordered tuple $(\lambda_{1},\cdots,\lambda_{n})$. Note that the number of labels is one less than the size of the underlying CNAT.
@@
Now suppose that we have a labelled CNAT $\mathcal{T}=(T,\lambda)$ such that the only internal dot in the top row of $T$ is the root of the tree. Such a tree must necessarily have a dot in the second row of the first column, which we call $\rho^{\prime}$. We define the top-row deletion of $\mathcal{T}$ as the labelled tree $\mathcal{T}^{\prime}:=(T^{\prime},\lambda^{\prime})$ where $T^{\prime}$ is the CNAT $T$ with top row deleted (i.e. the subtree rooted at $\rho^{\prime}$), and $\lambda^{\prime}:=\lambda\setminus\{\lambda_{1}\}$ is the set of labels $\lambda$ with its minimum element removed. This operation is shown on Figure 6, with the labelled CNAT on the left, and its top-row deleted labelled CNAT on the right.
@@
A similar decomposition was given for so-called alternative tableaux by Nadeau [22, Section 2]. Alternative tableaux are fillings of Ferrers diagrams with left- and up-arrows, and are known to be in bijection with tree-like tableaux (see [5]). Our decomposition places the construction directly in the CNAT setting, and is very natural in terms of the underlying tree structure.
We are now equipped to define the desired bijection from upper-diagonal CNATs to permutations. In fact, given any label set $\lambda$, we define a bijection from $\lambda$-labelled CNATs to set-permutations of the label set $\lambda$ ($\lambda$-permutations for short). Since upper-diagonal CNATs of size $n$ can be viewed as $[n-1]$-labelled CNATs, this gives the desired bijection.
Fix a label set $\lambda=(\lambda_{1},\cdots,\lambda_{n})$, and a $\lambda$-labelled CNAT $\mathcal{T}=(T,\lambda)$. We define a set-permutation $\Psi(\mathcal{T})$ of the label set $\lambda$ recursively as follows.
@@
Note that in Example 3.4 above, the left-to-right minima of the $\lambda$-permutation $\pi:=\Psi(\mathcal{T})$ are exactly the labels of the top-row dots of $\mathcal{T}$, i.e. $2,5,9$. Moreover, $\pi$ has $4=5-1$ descents, and there are $5$ empty rows, i.e. rows whose only dot is their leaf dot, in the underlying CNAT $T$ (the rows whose leaves are in columns labelled $2$, $4$, $5$, $9$, and $10$). It turns out that both of these observations are true in general, which is the main result of this section.
@@
In particular, if the label set is $\lambda=[n-1]$ for some $n\geq 1$, this gives us a bijection between upper-diagonal CNATs of size $n$ and $(n-1)$-permutations, providing a direct proof of Theorem 4.12 in [11], as announced.
@@
Recall that upper-diagonal CNATs are those whose associated permutation is $\mathrm{dec}_{[}n]=n(n-1)\cdots 1$. The corresponding permutation graph $G_{\mathrm{dec}_{[}n]}$ is simply the complete graph on $n$ vertices $K_{n}$. It is straightforward to see that acyclic orientations of $K_{n}$ are simply linear orders of its vertex set, i.e. $n$-permutations, and fixing the unique sink of these yields $(n-1)$-permutations, so that $t_{K_{n}}=(n-1)!$ by Proposition 2.2.
On the other hand, the proof of Theorem 2.12 in [13] goes through an alternate interpretation of the acyclic orientation number $T_{G_{\pi}}$ in terms of spanning trees of $G_{\pi}$ with no external activity (these are shown to correspond bijectively to CNATs with associated permutation $\pi$). In this context, Theorem 3.5 can be viewed as a bijection between spanning trees with no external activity and sink-rooted acyclic orientations in the special case of complete graphs. Such bijections do in fact exist in the general case (see e.g. [6] or [16]). The novelty of our bijection is placing this work directly in the CNAT setting, as well as the preservation of relevant statistics.
@@
Since top-row decomposition partitions the label set $\lambda$ into two (smaller) label sets, and top-row deletion removes the minimum label $\lambda_{1}$, we have that $\Psi(\mathcal{T})$ is a $\lambda$-permutation of $\lambda$ by a straightforward induction on the size of $\lambda$ using the recursive definition.
Let us now show that if there is a dot in the top row in the column labelled $\lambda_{i}$ of $\mathcal{T}$, then $\lambda_{i}$ is a left-to-right minimum of the permutation $\Psi(\mathcal{T})$. Again, we proceed by induction on $n$, the size of the label set $\lambda$. For $n=0$ the result is trivial. Now fix some $n>0$, and suppose we have shown this for all label sets $\lambda^{\prime}$ of size at most $n-1$. Fix a label set $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ of size $n$, and a $\lambda$-labelled CNAT $\mathcal{T}$.
@@
It thus remains to consider the case where $\mathcal{T}$ has at least two internal dots in the top row. In this case write the top-row decomposition $\mathcal{T}=\mathcal{T}^{\ell}\oplus\mathcal{T}^{r}$ of $\mathcal{T}$. By construction the labels of top-row internal dots of $\mathcal{T}$ are exactly those of top-row dots of $\mathcal{T}^{\ell}$, plus the label $\lambda^{r}_{1}$ of the right-most dot in that row (the root of $T^{r}$). By the induction hypothesis, $\Psi(\mathcal{T}^{\ell})$ is a permutation of $\lambda^{\ell}$ whose left-to-right minima are the column labels of the top-row dots of $\mathcal{T}^{\ell}$. In particular, this implies that the first letter in $\Psi(\mathcal{T}^{\ell})$ is strictly less than $\lambda^{r}_{1}$. Moreover, by the above argument $\Psi(\mathcal{T}^{r})$ is a permutation of $\lambda^{r}$ starting with $\lambda^{r}_{1}$ (the minimal element of $\lambda^{r}$). This implies that the left-to-right minima in the permutation $\Psi(\mathcal{T})=\Psi(\mathcal{T}^{r})\!\cdot\!\Psi(\mathcal{T}^{\ell})$ are exactly $\lambda^{r}_{1}$ and the left-to-right minima of $\Psi(\mathcal{T}^{\ell})$, and the result follows by applying the induction hypothesis to $\Psi(\mathcal{T}^{\ell})$.
To show that $\Psi$ is a bijection, we construct its inverse recursively, and show it maps left-to-right minima to top-row internal dots in the labelled CNAT. Let $\pi=\pi_{1}\cdots\pi_{n}$ be a permutation of a label set $\lambda=(\lambda_{1},\cdots,\lambda_{n})$. We define a $\lambda$-labelled CNAT $\mathcal{T}:=\Phi(\pi)$ as follows.
[…]
  • If $\pi_{1}=\lambda_{1}$ (the first element of $\pi$ is the minimum label in $\lambda$), then we first calculate recursively $\mathcal{T}^{\prime}=\Phi(\pi_{2}\cdots\pi_{n})$, and then let $\mathcal{T}$ be the unique $\lambda$-labelled CNAT whose top-row deletion is $\mathcal{T}^{\prime}$ (see Remark 3.3). By construction the CNAT has a unique dot in the top-row, which is in column labelled $\lambda_{1}$, giving the desired result.
@@
It remains to show that the number of empty rows in $\mathcal{T}$ is one plus the number of descents of $\Psi(\mathcal{T})$. For a labelled CNAT $\mathcal{T}$, resp. a permutation $\pi$, let $\mathrm{EmptyRow}\left([\right)\mathcal{T}]$, resp. $\mathrm{Desc}\left([\right)\pi]$, be its number of empty rows, resp. descents. Again, we proceed by induction on the size $n$ of the label set $\lambda$. The base case $n=0$ is trivial. For $n\geq 1$, suppose we have shown that $\mathrm{EmptyRow}\left([\right)\mathcal{T}^{\prime}]=1+\mathrm{Desc}\left([\right)\Psi(\mathcal{T}^{\prime})]$ for all $\lambda^{\prime}$ of size at most $n-1$ and $\lambda^{\prime}$-labelled CNATs $\mathcal{T}^{\prime}$.
Fix some label set $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ and $\lambda$-labelled CNAT $\mathcal{T}$.

### v4

@@ 2.3. Permutations, permutation patterns, and permutation graphs
For $n\geq 1$, we let $S_{n}$ be the set of permutations of size $n$ (called $n$-permutations for short). We will usually use standard one-line notation for permutations. That is, a permutation $\pi\in S_{n}$ is a word $\pi=\pi_{1}\cdots\pi_{n}$ on the alphabet $[n]$ such that every letter in $[n]$ appears exactly once in $\pi$. By convention, the unique 0-permutation is simply the empty word. For $i,j\in[n],i\neq j$, we write $i\prec_{\pi}j$ if $i$ appears before (to the left of) $j$ in the permutation (word) $\pi$.
Another useful way of viewing permutations is their plot. In this view, an $n$-permutation is a set of dots in an $n\times n$ grid such that every row and every column of the grid contains exactly one dot. For a permutation $\pi=\pi_{1}\cdots\pi_{n}$, we construct its plot by putting dots in column $i$ and row $\pi_{i}$ for $i=1,\ldots,n$. When connecting permutations to CNATs, it will be convenient to label columns from left-to-right and rows from top-to-bottom in this representation, as in Figure 2 below.
$112233445566$
@@
A permutation $\pi=\pi_{1}\cdots\pi_{n}$ is said to be reducible if there exists $1\leq k<n$ such that $\pi_{1}\cdots\pi_{k}$ is a $k$-permutation. A permutation is irreducible if it is not reducible. Such an index $k$ is called a breakpoint of the permutation $\pi$. For example, the permutation $\textbf{312}564$ is reducible, while $561243$ is irreducible. [^1] The terminology of (in)decomposable, or sum (in)decomposable, is also used sometimes in the literature instead of (ir)reducible. A fixed point of a permutation $\pi$ is an index $j\in[n]$ such that $\pi_{j}=j$. A left-to-right minimum is a letter $\pi_{i}$ such that $\pi_{j}>\pi_{i}$ for all $j<i$. For example, the permutation $\textbf{5}6\textbf{1}243$ has two left-to-right minima $5$ and $1$.
Given a permutation $\pi$ and two letters $i,j\in[n]$, we say that $(i,j)$ is an inversion of $\pi$ if $i>j$ and $i\prec_{\pi}j$. A descent of $\pi$ is an inversion consisting of consecutive letters in the word, that is $i=\pi_{k}$ and $j=\pi_{k+1}$ for some $k\in[n-1]$. For example, the permutation $\pi=561243$ has the following inversions (descents in bold): $(5,1),(5,2),(5,4),(5,3),\textbf{(6,1)},(6,2),(6,4),(6,3),\textbf{(4,3)}$. The permutation graph of $\pi$ is the graph with vertex set $[n]$ and edge set the set of inversions of $\pi$. We denote this graph $G_{\pi}$. Figure 3 shows the permutation graph of $\pi=561243$. Note that the labels of the vertices are the row labels in the plot of $\pi$.
112233445566112233445566334411225566
@@
This is a classical result on permutation graphs, see e.g. [18, Lemma 3.2]. From now on, we assume that all permutations are irreducible unless explicitly stated otherwise. One permutation that will play an important role in our paper, particularly in Section 3 and Theorem 4.23, is the decreasing permutation, defined by $\mathrm{dec}_{[}n]:=n(n-1)\cdots 1$ (for some $n\geq 1$). In $\mathrm{dec}_{[}n]$ every pair is an inversion, so the corresponding graph is the complete graph $K_{n}$.
Let $\pi$ be an $n$-permutation, and $\tau$ a $k$-permutation for some $k\leq n$. We say that $\pi$ contains the pattern $\tau$ if there exist indices $i_{1}<i_{2}<\cdots<i_{k}$ such that $\pi_{i_{1}},\pi_{i_{2}},\ldots,\pi_{i_{k}}$ appear in the same relative order as $\tau$. In other words, if we delete all columns and rows other than $(i_{1},\pi_{i_{1}}),\ldots,(i_{k},\pi_{i_{k}})$ from the plot of $\pi$, we get the plot of $\tau$. In this case, we call $\pi_{i_{1}}\cdots\pi_{i_{k}}$ an occurrence of the pattern $\tau$. If $\pi$ does not contain the pattern $\tau$, we say that $\pi$ avoids $\tau$. For example, the permutation $\pi=561243$ contains two occurrences of the pattern $321$, given by the bolded letters: $\textbf{5}612\textbf{43}$, $5\textbf{6}12\textbf{43}$. However, it avoids the pattern $4321$ since there is no sequence of $4$ letters in decreasing order.
@@
Patterns of a permutation $\pi$ are linked to induced subgraphs of its permutation graph $G_{\pi}$ via the following observation. Let $\pi$ be an $n$-permutation, and $\tau$ a $k$-permutation, for some $k\leq n$. Suppose that $\pi_{1},\ldots,\pi_{k}$ is an occurrence of the pattern $\tau$ in $\pi$. Then the induced subgraph $G\left[\{\pi_{1},\ldots,\pi_{k}\}\right]$ is isomorphic to the permutation graph $G_{\tau}$. Conversely, if there exist vertices $\pi_{1},\ldots,\pi_{k}$ such that $G\left[\{\pi_{1},\ldots,\pi_{k}\}\right]$ is isomorphic to $G_{\tau}$, then $\pi_{1},\ldots,\pi_{k}$ is an occurrence of some pattern $\tau^{\prime}$ in $\pi$, such that the graphs $G_{\tau}$ and $G_{\tau^{\prime}}$ are isomorphic. This yields the following.
@@ 2.4. Complete non-ambiguous trees and associated permutations
Following tree terminology, for a NAT $T$, we call the dot lying in the top-left cell the root dot, or simply the root, of $T$. Similarly, a leaf dot is a dot with no dots below it in the same column or to its right in the same row. An internal dot is a dot which is not a leaf dot (this includes the root, unless the NAT is a single dotted cell). Given a NAT, it will be convenient to label the columns $1,2,\dots,n$ from left to right, and the rows $1,2,\dots,m$ from top to bottom.
@@
It is straightforward to check that in a CNAT, every row and every column must have exactly one leaf dot (see e.g. [13, Section 4.3] or [11, Section 2.1]). The leaf dot in a column, resp. row, is simply its bottom-most, resp. right-most, dot. As such, the set of leaf dots of a CNAT $T$ of size $n$ forms the plot of an $n$-permutation $\pi$, which must be irreducible (see e.g. [11, Theorem 3.3]). We say that $\pi$ is the permutation associated with the CNAT $T$. For example, the CNAT on the left of Figure 4 has associated permutation $\pi=45312$. We say that a row or column of a CNAT is empty if it contains no dots other than its leaf dot.
Definition 2.12.
For a permutation $\pi$, we define the set $\mathrm{CNAT}\left([\right)\pi]$ to be the set of CNATs whose associated permutation is $\pi$, and $\mathrm{cnat}\left([\right)\pi]:=|\mathrm{CNAT}\left([\right)\pi]|$ to be the number of such CNATs.
2.5. Connecting CNATs to the acyclic orientation numbers
We are now ready to state the connection between CNATs and the acyclic orientation number. This is essentially a re-phrasing of [13, Corollary 33] if one uses the bijection betwen minimal recurrent configurations of the Abelian sandpile model and $s$-rooted acyclic orientations of the underlying graph (see [22, Corollary 3]).
@@ 2.5. Connecting CNATs to the acyclic orientation numbers
$\mathrm{cnat}\left([\right)\pi]=a_{G_{\pi}}.$
@@ 3. Upper-diagonal CNATs and permutations
In [11], the authors showed that for the decreasing permutation $\mathrm{dec}_{[}n]=n(n-1)\cdots 1$, we have $\mathrm{cnat}\left([\right){\mathrm{dec}_{[}n]}]=(n-1)!$. Their proof is via so-called tiered trees, with the authors exhibiting a bijection between upper-diagonal fully tiered trees on $n$ vertices and permutations of [n1][n-1]. In this section, we give a new proof of this result by exhibiting a direct bijection between the set $\mathrm{CNAT}\left([\right){\mathrm{dec}_{[}n]}]$ and the set of permutations of [n1][n-1]. Our bijection also maps certain statistics of CNATs to statistics on the corresponding permutation. Following notation from [11], we call a CNAT upper-diagonal if its associated permutation is a decreasing permutation.
@@
A labelled CNAT is a pair $\mathcal{T}=(T,\lambda)$ where $T$ is an upper-diagonal CNAT of size $(n+1)$ for some $n\geq 0$ and $\lambda=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{n}\}\subset\mathbb{N}^{n}$ is a set of $n$ distinct natural numbers. We call $\lambda$ the labels of $\mathcal{T}$, and say that $\mathcal{T}$ is a $\lambda$-labelled CNAT.
In the above, we think of $\lambda$ as labelling the columns of the CNAT $T$ other than the right-most column, as in Figures 5 and 6 below (the labels are written on top of the columns). With a slight abuse of notation, we assume that $\lambda_{1}<\lambda_{2}<\cdots<\lambda_{n}$, and use $\lambda$ to denote both the set of labels and the ordered tuple $(\lambda_{1},\ldots,\lambda_{n})$. Note that the number of labels is one less than the size of the underlying CNAT.
@@
Now suppose that we have a labelled CNAT $\mathcal{T}=(T,\lambda)$ such that the only internal dot in the top row of $T$ is the root of the tree. Such a tree must necessarily have a dot in the second row of the first column, which we call $\rho^{\prime}$. We define the top-row deletion of $\mathcal{T}$ as the labelled tree $\mathcal{T}^{\prime}:=(T^{\prime},\lambda^{\prime})$ where $T^{\prime}$ is the CNAT $T$ with top row deleted (i.e. the subtree rooted at $\rho^{\prime}$), and $\lambda^{\prime}:=\lambda\setminus\{\lambda_{1}\}$ is the set of labels $\lambda$ with its minimum element removed. This operation is shown in Figure 6, with the labelled CNAT on the left, and its top-row deleted labelled CNAT on the right.
@@
We are now equipped to define the desired bijection from upper-diagonal CNATs to permutations. In fact, given any label set $\lambda$, we define a bijection from $\lambda$-labelled CNATs to set-permutations of the label set $\lambda$ ($\lambda$-permutations for short). Since upper-diagonal CNATs of size $n$ can be viewed as [n1][n-1]-labelled CNATs, this gives the desired bijection.
Fix a label set $\lambda=(\lambda_{1},\ldots,\lambda_{n})$, and a $\lambda$-labelled CNAT $\mathcal{T}=(T,\lambda)$. We define a set-permutation $\Psi(\mathcal{T})$ of the label set $\lambda$ recursively as follows.
@@
Remark 3.4.
A similar decomposition was given for so-called alternative tableaux by Nadeau [21, Section 2]. Alternative tableaux are fillings of Ferrers diagrams with left- and up-arrows, and are known to be in bijection with tree-like tableaux [5], which are themselves in bijection with upper-diagonal CNATs [11].
More precisely, Nadeau’s decomposition uses a concept of free columns/rows for alternative tableaux. In the (labelled) CNAT setting, free columns correspond to the labels of dots in the top row, and Nadeau’s decomposition along free columns is essentially equivalent to our top-row decomposition operation. On the other hand, our top row deletion operation appears new, especially the relabelling that is required. In particular, this operation allows us to only act on the columns of the upper-diagonal CNAT, without needing to introduce equivalent decomposition operation(s) on its rows. Our construction also positions itself directly in the CNAT setting, and is very natural in terms of the underlying tree structure.
@@
Note that in Example 3.5 above, the left-to-right minima of the $\lambda$-permutation $\pi:=\Psi(\mathcal{T})$ are exactly the labels of the top-row dots of $\mathcal{T}$, i.e. $2,5,9$. Moreover, $\pi$ has 4=514=5-1 descents, and there are $5$ empty rows, i.e. rows whose only dot is their leaf dot, in the underlying CNAT $T$ (the rows whose leaves are in columns labelled $2$, 44, 55, 99, and $10$). It turns out that both of these observations are true in general, which is the main result of this section.
@@
In particular, if the label set is $\lambda=[n-1]$ for some $n\geq 1$, this gives us a bijection between upper-diagonal CNATs of size $n$ and (n1)(n-1)-permutations, providing a direct proof of Theorem 4.12 in [11], as announced.
@@
Recall that upper-diagonal CNATs are those whose associated permutation is $\mathrm{dec}_{[}n]=n(n-1)\cdots 1$. The corresponding permutation graph $G_{\mathrm{dec}_{[}n]}$ is simply the complete graph on $n$ vertices $K_{n}$. It is straightforward to see that acyclic orientations of $K_{n}$ are simply linear orders of its vertex set, i.e. $n$-permutations, and fixing the unique sink of these yields (n1)(n-1)-permutations, so that $a_{K_{n}}=(n-1)!$ by Proposition 2.1.
On the other hand, the proof of Theorem 2.13 in [13] goes through an alternate interpretation of the acyclic orientation number $a_{G_{\pi}}$ in terms of spanning trees of $G_{\pi}$ with no external activity (these are shown to correspond bijectively to CNATs with associated permutation $\pi$). In this context, Theorem 3.6 can be viewed as a bijection between spanning trees with no external activity and sink-rooted acyclic orientations in the special case of complete graphs. Such bijections do in fact exist in the general case (see e.g. [6] or [15]). The novelty of our bijection is placing this work directly in the CNAT setting, as well as the preservation of relevant statistics.
@@
Since top-row decomposition partitions the label set $\lambda$ into two (smaller) label sets, and top-row deletion removes the minimum label $\lambda_{1}$, we have that $\Psi(\mathcal{T})$ is a $\lambda$-permutation by a straightforward induction on the size of $\lambda$ using the recursive definition.
Let us now show that if there is a dot in the top row in the column labelled $\lambda_{i}$ of $\mathcal{T}$, then $\lambda_{i}$ is a left-to-right minimum of the permutation $\Psi(\mathcal{T})$. Again, we proceed by induction on $n$, the size of the label set $\lambda$. For $n=0$ the result is trivial. Now fix some $n>0$, and suppose we have shown this for all label sets $\lambda^{\prime}$ of size at most n1n-1. Fix a label set $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ of size $n$, and a $\lambda$-labelled CNAT $\mathcal{T}$.
@@
It thus remains to consider the case where $\mathcal{T}$ has at least two internal dots in the top row. In this case write the top-row decomposition $\mathcal{T}=\mathcal{T}^{\ell}\oplus\mathcal{T}^{r}$ of $\mathcal{T}$. By construction the labels of top-row internal dots of $\mathcal{T}$ are exactly those of top-row internal dots of $\mathcal{T}^{\ell}$, plus the label $\lambda^{r}_{1}$ of the right-most (internal) dot in that row (the root of $T^{r}$). By the induction hypothesis, $\Psi(\mathcal{T}^{\ell})$ is a permutation of $\lambda^{\ell}$ whose left-to-right minima are the column labels of the top-row dots of $\mathcal{T}^{\ell}$. In particular, this implies that the first letter in $\Psi(\mathcal{T}^{\ell})$ is strictly less than $\lambda^{r}_{1}$. Moreover, by the above argument $\Psi(\mathcal{T}^{r})$ is a permutation of $\lambda^{r}$ starting with $\lambda^{r}_{1}$ (the minimal element of $\lambda^{r}$). This implies that the left-to-right minima in the permutation $\Psi(\mathcal{T})=\Psi(\mathcal{T}^{r})\!\cdot\!\Psi(\mathcal{T}^{\ell})$ are exactly $\lambda^{r}_{1}$ and the left-to-right minima of $\Psi(\mathcal{T}^{\ell})$, and the result follows by applying the induction hypothesis to $\Psi(\mathcal{T}^{\ell})$.
To show that $\Psi$ is a bijection, we construct its inverse recursively, and show it maps left-to-right minima to top-row internal dots in the labelled CNAT. Let $\pi=\pi_{1}\cdots\pi_{n}$ be a permutation of a label set $\lambda=(\lambda_{1},\ldots,\lambda_{n})$. We define a $\lambda$-labelled CNAT $\mathcal{T}:=\Phi(\pi)$ as follows.
  • If $\lambda=\emptyset$ ($\pi$ is the empty word in this case), we set $T$ to be the CNAT consisting of a single dot, and $\mathcal{T}=(T,\emptyset)$ to be the labelled CNAT (base case).
  • If $\pi_{1}=\lambda_{1}$ (the first element of $\pi$ is the minimum label in $\lambda$), then we first calculate recursively $\mathcal{T}^{\prime}=\Phi(\pi_{2}\cdots\pi_{n})$, and then let $\mathcal{T}$ be the unique $\lambda$-labelled CNAT whose top-row deletion is $\mathcal{T}^{\prime}$ (see Remark 3.3). By construction the CNAT has a unique internal dot in the top-row, which is in column labelled $\lambda_{1}$, giving the desired result.
@@
It remains to show that the number of empty rows in $\mathcal{T}$ is one plus the number of descents of $\Psi(\mathcal{T})$. For a labelled CNAT $\mathcal{T}$, resp. a permutation $\pi$, let $\mathrm{EmptyRow}\left([\right)\mathcal{T}]$, resp. $\mathrm{Desc}\left([\right)\pi]$, be its number of empty rows, resp. descents. Again, we proceed by induction on the size $n$ of the label set $\lambda$. The base case $n=0$ is trivial. For $n\geq 1$, suppose we have shown that $\mathrm{EmptyRow}\left([\right)\mathcal{T}^{\prime}]=1+\mathrm{Desc}\left([\right)\Psi(\mathcal{T}^{\prime})]$ for all $\lambda^{\prime}$ of size at most n1n-1 and $\lambda^{\prime}$-labelled CNATs $\mathcal{T}^{\prime}$.
Fix some label set $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ and $\lambda$-labelled CNAT $\mathcal{T}$.